OpenStudy (anonymous):

5u/5u+10 Simplify it

OpenStudy (anonymous):

Do you mean \[\frac{5u}{5u}+10 \] in which case the \(5u\) cancels and leaves you with \(10\)

OpenStudy (anonymous):

Oh sorry 11

OpenStudy (anonymous):

\[\frac{5u}{5u}+10 = \frac{1}{1}+10 = 11\]

OpenStudy (anonymous):